OpenStudy (anonymous):

One number is five more than another. Their sum is eleven. Find both numbers.

OpenStudy (saifoo.khan):

Let's suppose the number be "x". So one number = x The other number is FIVE MORE than the first number so: x + 5 And their SUM is 11. So finally, x + (x+5) = 11

OpenStudy (anonymous):

So what is the smaller and larger value?

OpenStudy (saifoo.khan):

Solve for x to get the smaller value.

OpenStudy (anonymous):

8 and 3

OpenStudy (anonymous):

One number is three less than another. Their sum is eleven. Find the numbers. smaller value = larger value =

OpenStudy (saifoo.khan):

Solve for x to get the smaller value.

OpenStudy (anonymous):

how exactly?

OpenStudy (saifoo.khan):

Add like terms.. Isolate "x"

OpenStudy (anonymous):

please show me

OpenStudy (saifoo.khan):

x + (x+5) = 11 2x + 5 = 11 2x = 11 - 5 2x = 6 Do you get these steps?

OpenStudy (anonymous):

what would i have to do next since i have to find two numbers?

OpenStudy (anonymous):

are you stupid? obviously 8 is larger than 3.

OpenStudy (anonymous):

Matthewenit i thik you are the one who is stupid... can u read?

OpenStudy (saifoo.khan):

now: 2x = 6 x = 6/2 x = 3 This means taht our first tem is 3. Agree?

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

Now our second term was "x+5" And the value of "x" is 3.